Science Report – December 20th


Today’s work (Sol 2) was all about setup and preparation of the
equipment necessary for transplant of cultivars into the GreenHab.
The first task in this process was to review the temperature data from
the night of Sol 1 and determine if turning on the heater had the
desired effect of keeping the climate acceptable for plant growth.
While the hab remained above freezing all night (recorded low of 48 F)
we determined it was likely that the gradient effect was preventing
the warm air from getting down to the level of the temperature sensor.
Therefore, we installed a box fan above the cooler to help increase
the air circulation and hopefully reduce this gradient.  We also
experienced a high of 108.7 F at 12:17.  We are now manually using
both the heater and cooler for the coming days to try and maintain a
relatively constant temperature moving forward.  We are postponing the
move of plant from the hab to the GreenHab until tomorrow to ensure
acceptable temperature variation througout the day.  During EVA, we
evaluated the systems that are currently in the GreenHab and prepared
the equipment for the introduction of the plants tomorrow.  We also
turned the cooling fan on during the middle of the day again to help
regulate and equalize the temperature.  When the first group came back
from their EVA, Curtis & Co. was also at the hab and we were able to
have a very informative discussion with them about the plans for the
aquaponics system.  They will be back later in the week with necessary
equipment to help assist with that setup process as necessary.  The
last major accomplishment for us was the germination of several
species of seeds that will be moved into the GreenHab tomorrow.  These
species included Green Oak Lettuce, Red Oak Lettuce, Radish, Pinto
Bean, Kidney Bean, Popcorn, Carrot, Spinach, Onion and a mystery crop
whose seeds were discovered in the pantry upon our arrival into the
hab. Yay Mystery Crop!


Today I installed the weather station on the roof of the HAB with the
help of Crew Engineer Geoffrey Andrews. The weather station on the top
of the HAB will provide a good vantage point so that the solar
radiation sensor will be unobstructed. This data will be used by the
GreenHab scientists in order to quantify solar radiation changes
throughout the day. I plan on moving this weather station to the
GreenHab eventually to compare solar radiation levels.

One of the time-lapse cameras was placed in the GreenHab today to
record the progress in there throughout the duration of the mission.
Tomorrow I plan to install a time-lapse camera on our EVA at our final
destination to monitor the landscape.

Max/Min: Outdoor Temp – 10 F – 51 F

GreenHab Temp – 46 F – 108 F

Wind – 11.9 mph, gust – 12.3 mph

Solar Rad. Max – 592.7 W/m^2

UV Index – 3

Outdoor Humidity – 12% – 45%

Mars Self-Sleep Study:
Anselm, the crew journalist and I Connor have decided to embark on a
change of our sleep patterns in order to gauge the effects and
application to sleep patterns on Mars. We have decided we need more
time for work and a changed sleep pattern may help with this. Most
people sleep by getting 6-8 hours at night and being awake for 16-18
hours during the day. Instead of this pattern, Anselm and I have
decided to reduce our nightly chunk of sleep to 3-4 hours and two one
hour naps during the day. This will increase our overall awake time
during the day to 19 hr. I anticipate being tired the first day but
then adjusting quickly to this pattern.

Hopefully if this works and we become much more productive, we can
recommend these types of patterns for future astronauts. The Martian
day is a little over 24 hours and this has proven to mess up humans’
sleep cycles in certain tests. We want to explore alternative sleep

11-2 am sleep

(7 hr awake)

9-10 am nap

(6 hr awake)

4-5pm nap

(6 hr awake)

Science Report – December 19th




A calibrated pH/EC sensor was located in the GreenHab.  Last night the temperature in the GreenHab was monitored and dropped to approximately 19F which would be detrimental to plant growth.  This prompted us to ensure the heater in the GreenHab was on and functioning.  The temperature was set to 75F and will be monitored throughout the night to ensure optimum plant growth conditions.  If the temperature stabilizes we will setup the move plant into the GreenHab tomorrow and begin new germination.  We also investigated the equipment that is available for use.  We have questions regarding the hydroponic systems available.  Depending on correspondences about hydroponic capabilities we will alter our research plans and potential growing areas.



Today, I installed the indoor temperature sensor inside the GreenHAB so that we can monitor the indoor temperature there. In addition, I prepped my time-lapse cameras for use in the field. I set their exposure time and time between frames. We found the geologic maps of the area here in the GreenHab and are ready for an exploratory EVA to look for interesting sites